(D2-D'2-3D+3D')z=xy+7


1
Expert's answer
2021-05-31T06:48:08-0400

Rewriting the given equation :

"(D-D')(D+D'-3)=xy+7"

comparing this with standard equation

"(D-m_1D'-\alpha_1)(D-m_2D'-\alpha_2)z=F(x,y)"


"m_1 = 1;\space\space\space\space\space\space\space\space m_2 = – 1
\\ \alpha_1 = 0; \space\space\space\space\space\space\space\space\space \alpha_2 = 3"


Complimentary function CF "=e^{\alpha_1x}\Phi_1(y+m_1x)+e^{\alpha_2 x}\Phi_2(y+m_2 x)"

"=e^{0x}\Phi_1(y+x)+e^{3 x}\Phi_2(y-x)"

"=\Phi_1(y+x)+e^{3 x}\Phi_2(y-x)"


Now Particular Integrals corresponding to "xy" and "7" are as follows:

P.I.1 "=\dfrac{1}{(D-D')(D+D'-3)}xy"

Taking "3D" common from the denominator

"=\dfrac{1}{-3D\bigg(1-\dfrac{D'}{D}\bigg)\bigg(1-\dfrac{D}{3}-\dfrac{D'}{3}\bigg)}xy"

"=\dfrac{-1}{3D}\bigg(1-\dfrac{D'}{D}\bigg)^{-1}.\bigg(1-\dfrac{D}{3}-\dfrac{D'}{3}\bigg)^{-1}xy"

Using Binomial expansion (1+x)-1 and solving we get

"=\dfrac{-1}{3}\bigg[\dfrac{1}{D}+\dfrac{1}{3}+\dfrac{2}{3}\dfrac{D}{D'}+\dfrac{1}{9}D+\dfrac{1}{3}D'+\dfrac{1}{9}DD'+\dfrac{D'}{D^2}+.......\bigg]xy"


On taking terms of D, D', DD' to the power 1 since in F(x, y), x and y are in power 1 only

P.I.1"=-\dfrac{1}{3}\bigg[\dfrac{1}{2}x^2y+\dfrac{1}{3}xy+\dfrac{1}{3}x^2+\dfrac{1}{9}y+\dfrac{1}{3}x+\dfrac{1}{9}+\dfrac{1}{6}x^3\bigg]"


P.I.2"=\dfrac{-1}{3D}\bigg(1-\dfrac{D'}{D}\bigg)^{-1}.\bigg(1-\dfrac{D}{3}-\dfrac{D'}{3}\bigg)^{-1}7"

"=\dfrac{-7}{3}\bigg[\dfrac{1}{D}+\dfrac{1}{3}+\dfrac{2}{3}\dfrac{D}{D'}+\dfrac{1}{9}D+\dfrac{1}{3}D'+\dfrac{1}{9}DD'+\dfrac{D'}{D^2}+.......\bigg]"

Taking terms whose numerator does not have D or D' as they will be equal to zero

"=\dfrac{-7}{3}\dfrac{1}{D}-\dfrac{7}{9}"

"=\dfrac{7x}{3}-\dfrac{7}{9}"


Complete Solution, "z=C.F+P.I._1+P.I._2"

"z=\Phi_1(y+x)+e^{3 x}\Phi_2(y-x)-\dfrac{1}{3}\bigg[\dfrac{1}{2}x^2y+\dfrac{1}{3}xy+\dfrac{1}{3}x^2+\dfrac{1}{9}y+\dfrac{1}{3}x+\dfrac{1}{9}+\dfrac{1}{6}x^3\bigg]+\dfrac{7x}{3}-\dfrac{7}{9}"